Find the general solution of the following differential equations using method of undetermined coefficients: (i) y''−2y'+y =x2e5x,


1
Expert's answer
2022-03-01T15:49:26-0500

Corresponding homogeneous differential equation


"y''\u22122y'+y =0"

Characteristic (auxiliary) equation


"r^2-2r+1=0"

"r_1=r_2=1"

The general solution of the homogeneous differential equation is


"y_h=c_1e^x+c_2xe^x"

Find the particular solution of the non homogeneous differential equation


"y_p=(Ax^2+Bx+C)e^{5x}"

"y_p'=5(Ax^2+Bx+C)e^{5x}"

"+(2Ax+B)e^{5x}"

"y_p''=25(Ax^2+Bx+C)e^{5x}"

"+10(2Ax+B)e^{5x}+2Ae^{5x}"

Substitute


"25(Ax^2+Bx+C)e^{5x}"

"+10(2Ax+B)e^{5x}+2Ae^{5x}"

"-10(Ax^2+Bx+C)e^{5x}-2(2Ax+B)e^{5x}"

"+(Ax^2+Bx+C)e^{5x}=x^2e^{5x}"

"16Ax^2+(16B+16A)x+16C+8B+2A=x^2"

"A=1\/16"

"B=-1\/16"

"C=3\/128"

"y_p=(\dfrac{1}{16}x^2-\dfrac{1}{16}x+\dfrac{3}{128})e^{5x}"

The general solution of the non homogeneous differential equation is


"y=c_1e^x+c_2xe^x+(\dfrac{1}{16}x^2-\dfrac{1}{16}x+\dfrac{3}{128})e^{5x}"